OpenStudy (freckles):

Prove or disprove: Let A,B, and C be functions of x such that C(x)=A(x)+B(x). If C(x) is linear function, then A(x) and B(x) are also linear functions.

OpenStudy (perl):

easy to disprove this one

OpenStudy (freckles):

C(x) is linear but A(x)=sqrt(x) and B(x)=-sqrt(x) are not linear

OpenStudy (freckles):

Good job.

OpenStudy (perl):

counterexample: let A(x) = x^2 + 2x+3 , B(x) = -x^2 A(x) + B(x) = 2x +3 , which is linear

OpenStudy (freckles):

C(x)=0*

OpenStudy (freckles):

That example also works.